OpenStudy (turingtest):

Quick, who can prove x^0=1 Go!

OpenStudy (anonymous):

x/x = x^(1-1)=x^0=1

OpenStudy (turingtest):

three minutes, not bad, here's mine x^0=x^(0+0)=x^0(x^0) x^0=X^0/x^0=1

OpenStudy (anonymous):

technically you should justify division so it should be x^(0-0)=x^0/x^0.....not plus in the power....but hey, who's interested?